### v1

# S2ST-Omni: An Efficient and Scalable Multilingual Speech-to-Speech Translation Framework via Seamlessly Speech-Text Alignment and Streaming Speech Decoder

Abstract
Multilingual speech-to-speech translation (S2ST) aims to directly convert spoken utterances from multiple source languages into natural and intelligible speech in a target language. Despite recent progress, significant challenges remain: 1) achieving high-quality and low-latency S2ST remains a critical hurdle; 2) existing S2ST approaches heavily rely on large-scale parallel speech corpora, which are extremely difficult to collect. To address these issues, we propose S2ST-Omni, an efficient and scalable framework for multilingual speech-to-speech translation. To be specific, we decompose the S2ST task into speech-to-text translation (S2TT) and text-to-speech synthesis (TTS), unifying them within a single end-to-end speech-language model. To achieve high-quality S2TT while reducing dependence on parallel corpora, we leverage large-scale pretrained models—Whisper for audio understanding and Qwen 3.0 for text understanding. A lightweight speech adapter is further introduced to seamlessly align speech and text representations, enabling effective utilization of the pretrained multimodal knowledge. To ensure both translation quality and real-time performance, we adopt a pretrained streaming speech decoder in the TTS stage, generating target speech in an autoregressive manner. Extensive experiments on the CVSS benchmark demonstrate that S2ST-Omni outperforms state-of-the-art S2ST baselines while maintaining comparable latency, highlighting its effectiveness and practical potential for real-world deployment.
Index Terms—  Speech-to-speech translation, speech language model, speech-text alignment, streaming speech decoder
$\dagger$ denotes the corresponding author.
1 Introduction
Multilingual speech-to-speech translation (S2ST) seeks to directly convert spoken input from multiple source languages into speech in a target language. As a critical intersection of speech technology and cross-lingual understanding, it plays a fundamental role in enabling natural and efficient multilingual spoken communication [1, 2].
Conventional S2ST systems [3, 4, 5] predominantly adopt a statistically based cascaded architecture that decomposes the S2ST task into separate modules: automatic speech recognition (ASR) [6, 7], machine translation [8, 9], and text-to-speech synthesis (TTS) [10, 11]. Notable examples include the Verbmobil project [5], which employs a hybrid multi-engine approach to handle multilingual business dialogues, and the corpus-based framework proposed by [4], integrating multilingual speech recognition, text translation, and speech synthesis to achieve high-quality translation among English, Japanese, and Chinese. Nonetheless, cascaded pipelines inherently suffer from error propagation, rigid module interfaces, and limited real-time performance, restricting their effectiveness in complex practical applications [12].
With rapid advances in machine and deep learning, direct S2ST methods have emerged [13, 14, 12, 15], demonstrating better performance compared to traditional S2ST systems. TranSpeech [13] introduced a non-autoregressive (NAR) speech-to-speech translation model that addresses acoustic multimodality through a bilateral perturbation mechanism—combining style normalization and information enhancement to learn deterministic linguistic representations. Afterwards, DASpeech [14] used a two-pass NAR architecture that first generates target text using a directed acyclic graph-based linguistic decoder and then produces target speech with a FastSpeech 2 [16] acoustic decoder, showcasing the effectiveness and potential of unified models to capture cross-lingual speech mappings. Despite their impressive performance, these methods still rely heavily on large-scale, strictly aligned parallel speech corpora, limiting scalability to low-resource languages. Moreover, there remains substantial room for improving translation accuracy.
To overcome these problems, we propose S2ST-Omni, an end-to-end scalable speech language modeling framework for multilingual speech-to-speech translation. The core principle of our proposed S2ST-Omni lies in the efficient and seamless integration of large-scale multimodal pretrained models. This design significantly reduces reliance on parallel speech data while enabling efficient and high-quality S2ST. To elaborate, S2ST-Omni leverages powerful large-scale pretrained models: Whisper and for robust audio comprehension and Qwen 3.0 large language model (LLM) for deep text understanding. Meanwhile, a lightweight speech adapter is proposed to bridge the modality gap between speech and text, facilitating effective integration of multimodal pretrained knowledge.
Fig. 1: Overall architecture of the proposed S2ST-Omni framework. LSA denotes the lightweight speech adapter.
To further enhance translation quality while ensuring real-time usability, we employ a pretrained streaming speech decoder to synthesize the target utterance in a autoregressive (AR) fashion, making S2ST-Omni well-suited for practical deployment in real-world multilingual communication scenarios. Extensive experiments on the CVSS multilingual speech translation dataset demonstrate that S2ST-Omni achieves substantial improvements over current leading end-to-end and cascaded baselines, delivering higher translation quality alongside competitive latency and robustness, thereby illustrating its practical applicability and promising potential.
2 Related Work
This section briefly presents the background for the two-pass S2ST and speech language model for better legibility.
2.1 Two-pass S2ST
Due to the challenges of directly generating target speech [12], many research [17, 18] adopts a two-pass architecture, which decomposes the task into S2TT and TTS. While this paradigm benefits from intermediate textual supervision to enhance translation quality, it often relies heavily on large-scale triplet-parallel corpora (speech-text-speech), limiting its scalability in multilingual and low-resource scenarios. To address this limitation, ComSpeech [12] introduces a compositional framework that leverages existing S2TT and TTS models, reducing the dependence on triplet data. However, its reliance on a Vocabulary Adaptor, CTC decoding, and Phoneme Embedding results in a shallow representation at the phonetic level, lacking the capacity to model deep semantics or long-range contextual dependencies. As a consequence, it struggles to capture complex linguistic phenomena, particularly in semantically rich or ambiguous contexts.
2.2 Speech Language Model
Recent progressions in discrete audio representations [19, 20, 21] and self-supervised learning [22, 23] have led to powerful speech language models (SLMs), such as GSLM [24], VoxtLM [25], and VALL-E [26]. These models excel in natural speech generation and zero-shot synthesis. Inspired by their impressive performance, SLMs have been applied to S2ST by decoupling translation and acoustic unit generation, improving translation quality and flexibility [27, 28]. However, these approaches heavily depend on large-scale parallel speech corpora and struggle with effective speech-text modality fusion. In contrast, our proposed S2ST-Omni leverages pre-trained multimodal large-scale models with lightweight speech adapters to align multimodal speech and text representations, thereby reducing reliance on parallel triplet data while enhancing semantic understanding and enabling efficient end-to-end training.
3 METHODOLOGY
In this section, we delineate the detailed architecture of the proposed S2ST-Omni framework, whose overall schematic is illustrated in Fig. 1.
3.1 System Overview
As shown in Figure 1, our proposed S2ST-Omni framework comprises four key components: a pre-trained speech encoder, a lightweight speech adapter, a pre-trained LLM, and a pre-trained streaming speech decoder. Specifically, to achieve high-quality S2TT, we first leverage the pretrained Whisper model in conjunction with a lightweight speech adapter to effectively extract and align speech representations. These features are then transformed into representations compatible with the pretrained Qwen 3.0 LLM. The entire S2TT model is optimized using cross-entropy loss to ensure accurate and efficient speech understanding and recognition. For efficient and real-time speech synthesis, we integrate a pretrained streaming speech decoder that generates target audio in an AR manner.
3.2 S2TT Stage: Seamless Speech-Text Alignment
Benefiting from training on large-scale and weakly supervised multilingual corpora, the Whisper [29] architecture has demonstrated a strong generalization ability in capturing diverse acoustic and linguistic characteristics across a wide range of languages and conditions, including low-resource scenarios. These properties make it an ideal choice for multilingual speech understanding and translation tasks. Consequently, we adopt Whisper-medium as the speech encoder in our S2ST-Omni workflow.
Let $X_{s}\in\mathbb{R}^{T}$ denote the input source utterance, where $T$ is the duration of $X_{S}$, $\bm{T}\in\mathbb{R}^{L}$ represent its corresponding target transcript, where $L$ equals its text token length. The Whisper model first converts the raw waveform into a log-Mel spectrogram feature $F\in\mathbb{R}^{T_{1}\times D}$, where $T_{1}$ is the number of spectrogram frames and $D$ is the acoustic feature dimension, typically set to 80. Afterward, the log-Mel spectrogram $F$ is passed through the Whisper encoder, which consists of a stack of transformer blocks [30] comprising multi-head self-attention and feed-forward layers. The encoder also incorporates positional embeddings to retain temporal information and layer normalization for stable training. Through this architecture, the Whisper encoder is able to extract high-level speech representations that capture both local acoustic features and long-range contextual dependencies.
To fully leverage the powerful semantic understanding capabilities of the LLM, it is essential to bridge the representation gap between the Whisper encoder and the Qwen 3.0 LLM. To this end, we present a lightweight speech adapter $\mathcal{A}(\cdot)$ based on a multi-layer projection (MLP) module, which maps the encoded speech features into a compatible representation space compatible with the Qwen 3.0 LLM.
Formally, assume $\mathcal{F}_{\text{whisper}}\in\mathbb{R}^{T_{w}\times D_{w}}$ denote the encoded speech features produced by the Whisper encoder. These features are further transformed by a four-layer feed-forward speech adapter $\mathcal{A}(\cdot)$, yielding the adapted representation $\mathcal{F}_{\text{adapt}}$ as:
$\begin{split}\mathcal{H}_{0}&=\mathcal{F}_{\text{whisper}}\end{split}$ (1)
$\mathcal{H}_{i}=\sigma(\mathcal{H}_{i-1}\mathbf{W}_{i}+\mathbf{b}_{i}),\quad\forall i\in\{1,2,3,4\}$ (2)
$\begin{split}\mathcal{F}_{\text{adapt}}&=\mathcal{H}_{4},\end{split}$ (3)
where $\mathcal{F}_{\text{adapt}}\in\mathbb{R}^{T_{a}\times D_{a}}$, $\mathbf{W}_{i}$ and $\mathbf{b}_{i}$ are the learnable weight matrices and bias vectors of the $i$-th linear layer, respectively, and $\sigma(\cdot)$ denotes the LeakyReLU activation function with a negative slope of 0.1. The operator $\prod$ here represents nested functional composition. In this way, S2ST-Omni enables efficient alignment of speech features with the language model’s input space while maintaining computational simplicity and training stability. Finally, the adjusted representation $\mathcal{F}_{\text{adapt}}$ is input into the Qwen 3.0 LLM for S2TT, with the guidance of the conventional cross-entropy loss:
$\mathcal{L}_{\text{S2TT}}=-\sum_{l=1}^{L}\log P(t_{l}\mid t_{<l},\mathcal{F}_{\text{adapt}})$ (4)
where ${t_{l}}$ denotes the $l$-th predicted target text token, and $t_{<l}$ represents the autoregressively generated text tokens preceding the $l$-th position.
3.3 TTS Stage: Streaming Speech Decoder
To achieve a balance between synthesis latency and translation quality, we incorporate a pretrained streaming speech decoder, CosyVoice2 [31], into the TTS component of S2ST-Omni.
Table 1: ASR-BLEU score comparison among all evaluated S2ST methods.
[…]
ASR-BLEU $\uparrow$ ASR-BLEU $\uparrow$ ASR-BLEU $\uparrow$
S2UT [32] 22.23 2.99 18.53
Translatotron 2 [17] 26.12 16.92 22.98
DASpeech [14] 25.03 16.14 21.37
UnitY [18] 26.90 16.36 24.06
ComSpeech [12] 28.15 18.16 24.80
StreamSpeech [15] 28.45 20.93 27.25
S2ST-Omni ??? ??? ???
Specifically, we first employ a supervised speech tokenizer1 to convert the target waveform $X_{T}$ into a sequence of discrete speech tokens $Y=[y_{1},y_{2},\ldots,y_{M}]$, sampled at a rate of 25 tokens per second, where each token $y_{i}\in\{0,1,\ldots,6561\}$. The predicted text output from the Qwen 3.0 large language model (LLM), denoted as $\bm{\hat{T}}$, is subsequently fed into a text-to-speech language model, which autoregressively generates the predicted speech token sequence $\hat{\mathbf{Y}}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{M}]$. To enhance the quality of synthesized speech, the TTS model is optimized via a cross-entropy loss function defined as:
$\mathcal{L}_{\text{TTS}}=-\sum_{i=1}^{M}\log P(y_{i}\mid\hat{y}_{<i},\bm{\hat{T}}),$ (5)
where $P(y_{i}\mid\hat{y}_{<i},\bm{\hat{T}})$ represents the likelihood of correctly predicting the $i$-th token, conditioned on the previously generated tokens and the text embedding $\bm{\hat{T}}$ derived from the predicted text. Following this, the predicted speech tokens $\hat{\mathbf{Y}}$ are reconstructed into Mel-spectrogram features of the target utterance $\bm{\hat{\mathbf{F}}}_{\mathrm{mel}}$ using a chunk-aware, causal conditional flow matching (CFM) [33, 34] model. This model is conditioned on prompt speaker embeddings to ensure speaker-consistent speech synthesis. In particular, every $K$ generated speech tokens are grouped as a chunk, which serves as the input unit for Mel-spectrogram generation. In our case, we fine-tune the CFM model starting from the pretrained CosyVoice2-CFM checkpoint2, while retaining the original training objective to preserve model stability and synthesis fidelity. More implementation details and training configurations can be referred to [31]. Ultimately, the predicted target Mel-spectrogram $\bm{\hat{\mathbf{F}}}_{\mathrm{mel}}$ is passed through a pretrained HifiGAN Vocoder3 [35], which reconstructs the final waveform to produce the final target speech.
1https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B/blob/main/speech_tokenizer_v2.onnx2https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B/blob/main/flow.pt3https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B/blob/main/hift.pt
4 EXPERIMENTS
In this section, we first describe the datasets used for the S2ST task. We then present the implementation details of the proposed S2ST-Omni model and briefly introduce the baseline systems used for comparison. Finally, we report and analyze the overall experimental results.
4.1 Experimental Settings
4.1.1 Databases
All experiments are conducted on the CVSS corpus [17], a public large-scale multilingual S2ST benchmark. The CVSS corpus consists of aligned triplets in the format of <source speech, target text, target speech>  across 21 source languages translated into English.
[…]
4.1.2 Implementation Details
In all experiments, we employ the AdamW optimizer to train the proposed S2ST-Omni model, with an initial learning rate set to $1\times 10^{-5}$. The batch sizes for the speech-to-text translation (S2TT) and text-to-speech (TTS) stages are 32 and 128, respectively. All models are trained for 20 epochs using the PyTorch framework on four NVIDIA A100 GPUs. The system prompt for the Qwen 3.0 large language model (LLM) is defined as: “You are an expert in speech-to-speech translation, responsible for accurately translating multilingual input speech into corresponding English text.”
Regarding the evaluating metric, we employ the commonly used ASR-BLEU to assess performance. In practice, the translated speech is first transcribed into its corresponding transcript using a pre-trained ASR model4, after which the BLEU score [36] is computed using the SacreBLEU toolkit5 to assess translation quality and statistical significance.
4https://dl.fbaipublicfiles.com/fairseq/wav2vec/wav2vec_vox_960h_pl.pt5https://github.com/mjpost/sacrebleu
4.1.3 Baselines
To ensure a fair and comprehensive comparison, we evaluate our S2ST-Omni system against six well-established S2ST methods:
  • S2UT [32]: A single-stage NAR model that directly translates speech into discrete units.
  • Translatotron 2 [17]: An end-to-end, one-pass system for direct speech-to-speech translation.
  • DASpeech [14]: A two-stage non-autoregressive framework that first converts speech to text and subsequently from text to discrete units.
  • UnitY [18]: A fully AR pipeline employing two passes to improve translation accuracy.
  • ComSpeech [12]: A composite model integrating pretrained S2TT and TTS modules into a direct S2ST system via contrastive learning.
  • StreamSpeech [15]: A unified multi-task model that performs direct simultaneous S2ST by jointly learning translation and timing policies for real-time streaming input.
4.2 Results and Analysis
To validate the effectiveness of the proposed method, we compare the translation performance of S2ST-Omni against several state-of-the-art S2ST systems. A comprehensive summary of the results is presented in Table 1.
[…]
  • Leveraging large-scale pre-trained models: By integrating Whisper’s robust acoustic representation learning with the powerful semantic modeling capabilities of Qwen 3.0, S2ST-Omni effectively exploits the rich knowledge embedded in large-scale pre-training. This substantially reduces the reliance on extensive parallel S2ST data, which is often scarce and costly to obtain.
  • Seamless speech-text alignment: The introduction of lightweight speech adapters enables the effective mapping of Whisper’s speech representations into the language space of Qwen 3.0. This design bridges the modality gap between speech and text, facilitates efficient knowledge transfer, and enhances alignment quality for downstream translation tasks.
  • Streaming speech decoder for TTS: A streaming speech decoder is employed during the text-to-speech synthesis stage, achieving real-time, low-latency voice generation without compromising translation quality—an essential requirement for practical deployment.
In summary, the experimental results provide strong evidence for the effectiveness and practicality of S2ST-Omni in multilingual speech-to-speech translation.
5 CONCLUSIONS
In this study, we introduce S2ST-Omni, a unified and efficient framework for multilingual speech-to-speech translation. By decomposing the S2ST task into speech-to-text translation and text-to-speech synthesis, and integrating large-scale pre-trained models such as Whisper and Qwen 3.0 with a lightweight speech adapter, S2ST-Omni achieves high-quality translation while reducing reliance on parallel speech data. Furthermore, the adoption of a streaming speech decoder enables low-latency speech generation suitable for real-time applications. Extensive experiments on the CVSS benchmark confirm that S2ST-Omni achieves better translation quality than existing state-of-the-art systems, demonstrating its effectiveness and superiority in practical multilingual S2ST applications.

### v2

@@
-S2ST-Omni: An Efficient and Scalable Multilingual Speech-to-Speech Translation Framework via Seamlessly Speech-Text Alignment and Streaming Speech Decoder
+S2ST-Omni: An Efficient and Scalable Speech-to-Speech Translation Framework via Seamlessly Speech-Text Alignment and Streaming Speech Decoder
 Abstract
-Multilingual speech-to-speech translation (S2ST) aims to directly convert spoken utterances from multiple source languages into natural and intelligible speech in a target language.
-Despite recent progress, significant challenges remain: 1) achieving high-quality and low-latency S2ST remains a critical hurdle; 2) existing S2ST approaches heavily rely on large-scale parallel speech corpora, which are extremely difficult to collect.
-To address these issues, we propose S2ST-Omni, an efficient and scalable framework for multilingual speech-to-speech translation.
-To be specific, we decompose the S2ST task into speech-to-text translation (S2TT) and text-to-speech synthesis (TTS), unifying them within a single end-to-end speech-language model.
-To achieve high-quality S2TT while reducing dependence on parallel corpora, we leverage large-scale pretrained models—Whisper for audio understanding and Qwen 3.0 for text understanding.
-A lightweight speech adapter is further introduced to seamlessly align speech and text representations, enabling effective utilization of the pretrained multimodal knowledge.
-To ensure both translation quality and real-time performance, we adopt a pretrained streaming speech decoder in the TTS stage, generating target speech in an autoregressive manner.
-Extensive experiments on the CVSS benchmark demonstrate that S2ST-Omni outperforms state-of-the-art S2ST baselines while maintaining comparable latency, highlighting its effectiveness and practical potential for real-world deployment.
+Multilingual speech-to-speech translation (S2ST) aims to directly convert spoken utterances from multiple source languages into fluent and intelligible speech in a target language.
+Despite recent progress, several critical challenges persist: 1) achieving high-quality and low-latency S2ST remains a significant obstacle; 2) most existing S2ST methods rely heavily on large-scale parallel speech corpora, which are difficult and resource-intensive to obtain.
+To tackle these challenges, we introduce S2ST-Omni, a novel, efficient, and scalable framework tailored for multilingual speech-to-speech translation.
+To enable high-quality S2TT while mitigating reliance on large-scale parallel speech corpora, we leverage powerful pretrained models: Whisper for robust audio understanding and Qwen 3.0 for advanced text comprehension.
+A lightweight speech adapter is introduced to bridge the modality gap between speech and text representations, facilitating effective utilization of pretrained multimodal knowledge.
+To ensure both translation accuracy and real-time responsiveness, we adopt a streaming speech decoder in the TTS stage, which generates the target speech in an autoregressive manner.
+Extensive experiments conducted on the CVSS benchmark demonstrate that S2ST-Omni consistently surpasses several state-of-the-art S2ST baselines in translation quality, highlighting its effectiveness and superiority.
 Index Terms—  Speech-to-speech translation, speech language model, speech-text alignment, streaming speech decoder
@@
 StreamSpeech [15] 28.45 20.93 27.25
-S2ST-Omni ??? ??? ???
+S2ST-Omni 67.45 71.20 68.28
 Specifically, we first employ a supervised speech tokenizer1 to convert the target waveform $X_{T}$ into a sequence of discrete speech tokens $Y=[y_{1},y_{2},\ldots,y_{M}]$, sampled at a rate of 25 tokens per second, where each token $y_{i}\in\{0,1,\ldots,6561\}$.
@@
 The batch sizes for the speech-to-text translation (S2TT) and text-to-speech (TTS) stages are 32 and 128, respectively.
-All models are trained for 20 epochs using the PyTorch framework on four NVIDIA A100 GPUs.
-The system prompt for the Qwen 3.0 large language model (LLM) is defined as: “You are an expert in speech-to-speech translation, responsible for accurately translating multilingual input speech into corresponding English text.”
+All models are trained for 100,000 steps using the PyTorch framework on four NVIDIA A100 GPUs.
+The system prompt for the used Qwen 3.0 4B LLM is defined as: “You are an expert in speech-to-speech translation, responsible for accurately translating multilingual input speech into corresponding English text.”
 Regarding the evaluating metric, we employ the commonly used ASR-BLEU to assess performance.
 In practice, the translated speech is first transcribed into its corresponding transcript using a pre-trained ASR model4, after which the BLEU score [36] is computed using the SacreBLEU toolkit5 to assess translation quality and statistical significance.
-4https://dl.fbaipublicfiles.com/fairseq/wav2vec/wav2vec_vox_960h_pl.pt5https://github.com/mjpost/sacrebleu
+4https://dl.fbaipublicfiles.com/fairseq/wav2vec/wav2vec_vox_960h_pl.pt5https://github.com/mjpost/sacrebleu6https://facebookresearch.github.io/stopes/docs/eval/blaser
 4.1.3 Baselines
@@
 4.2 Results and Analysis
-To validate the effectiveness of the proposed method, we compare the translation performance of S2ST-Omni against several state-of-the-art S2ST systems.
-A comprehensive summary of the results is presented in Table 1.
-[…]
-[…]
+To validate the effectiveness of the proposed approach, we compare the translation performance of S2ST-Omni against six state-of-the-art S2ST systems.
+A comprehensive summary of the results is provided in Table 1.
+As shown in the table, S2ST-Omni consistently and significantly outperforms all baselines on the CVSS benchmark across the evaluated language pairs, with gains exceeding 38 BLEU points in certain cases.
+These results highlight the strong generalization capability of the proposed framework across linguistically diverse settings.
+Specifically, in the French $\rightarrow$ English direction, S2ST-Omni achieves an ASR-BLEU score of 67.45, markedly surpassing the strongest baseline (StreamSpeech, 28.45).
+This substantial improvement underscores the model’s robustness in handling phonetic variability and prosodic features characteristic of French.
+For German $\rightarrow$ English, the model attains a remarkable 71.20 ASR-BLEU score—the highest among all language pairs—demonstrating its effectiveness in managing syntactic complexity and compound structures inherent to German.
+In the Spanish $\rightarrow$ English setting, S2ST-Omni achieves a strong 68.28 ASR-BLEU score, again outperforming all baselines by a wide margin.
+This result reflects the model’s ability to capture the rich morphological patterns of Spanish and generate fluent, accurate translations.
+The outstanding performance of S2ST-Omni can be attributed to several key architectural innovations:
   • Leveraging large-scale pre-trained models: By integrating Whisper’s robust acoustic representation learning with the powerful semantic modeling capabilities of Qwen 3.0, S2ST-Omni effectively exploits the rich knowledge embedded in large-scale pre-training.
@@
   • Streaming speech decoder for TTS: A streaming speech decoder is employed during the text-to-speech synthesis stage, achieving real-time, low-latency voice generation without compromising translation quality—an essential requirement for practical deployment.
-In summary, the experimental results provide strong evidence for the effectiveness and practicality of S2ST-Omni in multilingual speech-to-speech translation.
+Overall, the substantial and consistent improvements across all language pairs validate the robustness and efficacy of the proposed S2ST-Omni framework, establishing it as a new state-of-the-art in direct speech-to-speech translation with promising potential for real-world multilingual applications.
 5 CONCLUSIONS

### v3

@@
-S2ST-Omni: An Efficient and Scalable Speech-to-Speech Translation Framework via Seamlessly Speech-Text Alignment and Streaming Speech Decoder
+S2ST-Omni: An Efficient and Scalable Multilingual Speech-to-Speech Translation Framework via Seamless Speech-Text Alignment and Streaming Speech Generation
 Abstract
 Multilingual speech-to-speech translation (S2ST) aims to directly convert spoken utterances from multiple source languages into fluent and intelligible speech in a target language.
-Despite recent progress, several critical challenges persist: 1) achieving high-quality and low-latency S2ST remains a significant obstacle; 2) most existing S2ST methods rely heavily on large-scale parallel speech corpora, which are difficult and resource-intensive to obtain.
+Despite recent progress, several critical challenges persist: 1) achieving high-quality S2ST remains a significant obstacle; 2) most existing S2ST methods rely heavily on large-scale parallel speech corpora, which are difficult and resource-intensive to obtain.
 To tackle these challenges, we introduce S2ST-Omni, a novel, efficient, and scalable framework tailored for multilingual speech-to-speech translation.
+Specifically, we decompose S2ST into speech-to-text translation (S2TT) and text-to-speech synthesis (TTS).
 To enable high-quality S2TT while mitigating reliance on large-scale parallel speech corpora, we leverage powerful pretrained models: Whisper for robust audio understanding and Qwen 3.0 for advanced text comprehension.
 A lightweight speech adapter is introduced to bridge the modality gap between speech and text representations, facilitating effective utilization of pretrained multimodal knowledge.
-To ensure both translation accuracy and real-time responsiveness, we adopt a streaming speech decoder in the TTS stage, which generates the target speech in an autoregressive manner.
+To ensure both translation accuracy and real-time responsiveness, we adopt a streaming speech generation model in the TTS stage, which generates the target speech in an autoregressive manner.
 Extensive experiments conducted on the CVSS benchmark demonstrate that S2ST-Omni consistently surpasses several state-of-the-art S2ST baselines in translation quality, highlighting its effectiveness and superiority.
@@
 […]
-ASR-BLEU $\uparrow$ ASR-BLEU $\uparrow$ ASR-BLEU $\uparrow$
-S2UT [32] 22.23 2.99 18.53
-Translatotron 2 [17] 26.12 16.92 22.98
-DASpeech [14] 25.03 16.14 21.37
-UnitY [18] 26.90 16.36 24.06
-ComSpeech [12] 28.15 18.16 24.80
-StreamSpeech [15] 28.45 20.93 27.25
-S2ST-Omni 67.45 71.20 68.28
+S2UT [32] 22.23 2.99 18.53 14.58
+Translatotron 2 [17] 26.12 16.92 22.98 22.01
+DASpeech [14] 25.03 16.14 21.37 20.85
+UnitY [18] 26.90 16.36 24.06 22.44
+ComSpeech [12] 28.15 18.16 24.80 23.70
+StreamSpeech [15] 28.45 20.93 27.25 25.54
+S2ST-Omni 67.45 71.20 68.28 68.98
 Specifically, we first employ a supervised speech tokenizer1 to convert the target waveform $X_{T}$ into a sequence of discrete speech tokens $Y=[y_{1},y_{2},\ldots,y_{M}]$, sampled at a rate of 25 tokens per second, where each token $y_{i}\in\{0,1,\ldots,6561\}$.
@@
 The system prompt for the used Qwen 3.0 4B LLM is defined as: “You are an expert in speech-to-speech translation, responsible for accurately translating multilingual input speech into corresponding English text.”
-Regarding the evaluating metric, we employ the commonly used ASR-BLEU to assess performance.
-In practice, the translated speech is first transcribed into its corresponding transcript using a pre-trained ASR model4, after which the BLEU score [36] is computed using the SacreBLEU toolkit5 to assess translation quality and statistical significance.
+Regarding evaluation, we adopt two widely used metrics: ASR-BLEU and BLASER 2.0, to comprehensively assess the performance of all comparative methods.
+Specifically, for ASR-BLEU, the translated speech is first transcribed into text using a pretrained ASR model4.
+The resulting transcripts are then evaluated using the BLEU score [36], calculated with the SacreBLEU toolkit5, to quantify translation quality and statistical significance.
+For BLASER, we follow [15], employing the BLASER 2.0 toolkit6 to measure cross-lingual semantic similarity between source and translated utterances.
 4https://dl.fbaipublicfiles.com/fairseq/wav2vec/wav2vec_vox_960h_pl.pt5https://github.com/mjpost/sacrebleu6https://facebookresearch.github.io/stopes/docs/eval/blaser
@@
   • StreamSpeech [15]: A unified multi-task model that performs direct simultaneous S2ST by jointly learning translation and timing policies for real-time streaming input.
+Table 2: Offline S2ST performance of StreamSpeech, evaluated with BLASER 2.0.
+[…]
+BLASER 2.0 (Unsupervised)
+UnitY 0.4467 0.4473 0.5090 0.5116
+StreamSpeech 0.4486 0.4491 0.5155 0.5178
+S2ST-Omni 0.6523 0.7138 0.7262 0.6974
+BLASER 2.0 (QE)
+UnitY 3.1674 3.1772 3.3020 3.3278
+StreamSpeech 3.1779 3.1872 3.3442 3.3669
+S2ST-Omni 3.6475 3.9284 3.9648 3.8469
+BLASER 2.0 (Ref)
+UnitY 3.1744 3.1965 3.2213 3.2638
+StreamSpeech 3.1989 3.2200 3.3146 3.3525
+S2ST-Omni 3.8925 4.0897 4.0634 4.0152
 4.2 Results and Analysis
 To validate the effectiveness of the proposed approach, we compare the translation performance of S2ST-Omni against six state-of-the-art S2ST systems.
-A comprehensive summary of the results is provided in Table 1.
-As shown in the table, S2ST-Omni consistently and significantly outperforms all baselines on the CVSS benchmark across the evaluated language pairs, with gains exceeding 38 BLEU points in certain cases.
+A comprehensive summary of the results is provided in the Table 1 and Table 2.
+As shown in Table 1, S2ST-Omni consistently and significantly outperforms all baselines on the CVSS benchmark across the evaluated language pairs, with ASR-BLEU improvements exceeding in certain cases.
 These results highlight the strong generalization capability of the proposed framework across linguistically diverse settings.
@@
 This result reflects the model’s ability to capture the rich morphological patterns of Spanish and generate fluent, accurate translations.
-The outstanding performance of S2ST-Omni can be attributed to several key architectural innovations:
-  • Leveraging large-scale pre-trained models: By integrating Whisper’s robust acoustic representation learning with the powerful semantic modeling capabilities of Qwen 3.0, S2ST-Omni effectively exploits the rich knowledge embedded in large-scale pre-training.
-This substantially reduces the reliance on extensive parallel S2ST data, which is often scarce and costly to obtain.
-  • Seamless speech-text alignment: The introduction of lightweight speech adapters enables the effective mapping of Whisper’s speech representations into the language space of Qwen 3.0.
-This design bridges the modality gap between speech and text, facilitates efficient knowledge transfer, and enhances alignment quality for downstream translation tasks.
-  • Streaming speech decoder for TTS: A streaming speech decoder is employed during the text-to-speech synthesis stage, achieving real-time, low-latency voice generation without compromising translation quality—an essential requirement for practical deployment.
+Furthermore, to directly evaluate the quality of the converted speech, we also conduct BLASER comparisons which contain Unsupervised (score from 0 to 1), QE (score from 1 to 5),and Ref (score from 1 to 5) modes.
+As illustrated in Table 2, we can easily observe that the proposed S2ST-Omni system consistently achieves the best performance across all evaluation modes.
+In particular, S2ST-Omni achieves 34.7%, 14.3% and 19.8% relative BLASER improvements over the secondary StreamSpeech in terms of Unsupervised, QE, and Ref modes, highlighting the effectiveness and superiority of the proposed approach.
 Overall, the substantial and consistent improvements across all language pairs validate the robustness and efficacy of the proposed S2ST-Omni framework, establishing it as a new state-of-the-art in direct speech-to-speech translation with promising potential for real-world multilingual applications.
@@
 By decomposing the S2ST task into speech-to-text translation and text-to-speech synthesis, and integrating large-scale pre-trained models such as Whisper and Qwen 3.0 with a lightweight speech adapter, S2ST-Omni achieves high-quality translation while reducing reliance on parallel speech data.
-Furthermore, the adoption of a streaming speech decoder enables low-latency speech generation suitable for real-time applications.
-Extensive experiments on the CVSS benchmark confirm that S2ST-Omni achieves better translation quality than existing state-of-the-art systems, demonstrating its effectiveness and superiority in practical multilingual S2ST applications.
+Furthermore, the adoption of a streaming speech generation model enables low-latency speech generation suitable for real-time applications.
+Extensive experiments on the CVSS benchmark confirm that S2ST-Omni achieves better translation quality than existing state-of-the-art systems, demonstrating its effectiveness and superiority.

### v4

@@
 StreamSpeech [15] 28.45 20.93 27.25 25.54
-S2ST-Omni 67.45 71.20 68.28 68.98
+S2ST-Omni ??? ??? ??? ???
 Specifically, we first employ a supervised speech tokenizer1 to convert the target waveform $X_{T}$ into a sequence of discrete speech tokens $Y=[y_{1},y_{2},\ldots,y_{M}]$, sampled at a rate of 25 tokens per second, where each token $y_{i}\in\{0,1,\ldots,6561\}$.
@@
 StreamSpeech 0.4486 0.4491 0.5155 0.5178
-S2ST-Omni 0.6523 0.7138 0.7262 0.6974
+S2ST-Omni ??? ??? ??? ???
 BLASER 2.0 (QE)
@@
 StreamSpeech 3.1779 3.1872 3.3442 3.3669
-S2ST-Omni 3.6475 3.9284 3.9648 3.8469
+S2ST-Omni ??? ??? ??? ???
 BLASER 2.0 (Ref)
@@
 StreamSpeech 3.1989 3.2200 3.3146 3.3525
-S2ST-Omni 3.8925 4.0897 4.0634 4.0152
+S2ST-Omni ??? ??? ??? ???
 4.2 Results and Analysis
@@
 A comprehensive summary of the results is provided in the Table 1 and Table 2.
-As shown in Table 1, S2ST-Omni consistently and significantly outperforms all baselines on the CVSS benchmark across the evaluated language pairs, with ASR-BLEU improvements exceeding in certain cases.
-These results highlight the strong generalization capability of the proposed framework across linguistically diverse settings.
-Specifically, in the French $\rightarrow$ English direction, S2ST-Omni achieves an ASR-BLEU score of 67.45, markedly surpassing the strongest baseline (StreamSpeech, 28.45).
-This substantial improvement underscores the model’s robustness in handling phonetic variability and prosodic features characteristic of French.
-For German $\rightarrow$ English, the model attains a remarkable 71.20 ASR-BLEU score—the highest among all language pairs—demonstrating its effectiveness in managing syntactic complexity and compound structures inherent to German.
-In the Spanish $\rightarrow$ English setting, S2ST-Omni achieves a strong 68.28 ASR-BLEU score, again outperforming all baselines by a wide margin.
-This result reflects the model’s ability to capture the rich morphological patterns of Spanish and generate fluent, accurate translations.
-Furthermore, to directly evaluate the quality of the converted speech, we also conduct BLASER comparisons which contain Unsupervised (score from 0 to 1), QE (score from 1 to 5),and Ref (score from 1 to 5) modes.
-As illustrated in Table 2, we can easily observe that the proposed S2ST-Omni system consistently achieves the best performance across all evaluation modes.
-In particular, S2ST-Omni achieves 34.7%, 14.3% and 19.8% relative BLASER improvements over the secondary StreamSpeech in terms of Unsupervised, QE, and Ref modes, highlighting the effectiveness and superiority of the proposed approach.
-Overall, the substantial and consistent improvements across all language pairs validate the robustness and efficacy of the proposed S2ST-Omni framework, establishing it as a new state-of-the-art in direct speech-to-speech translation with promising potential for real-world multilingual applications.
 5 CONCLUSIONS

### v5

@@
-S2ST-Omni: An Efficient and Scalable Multilingual Speech-to-Speech Translation Framework via Seamless Speech-Text Alignment and Streaming Speech Generation
+S2ST-Omni: An Efficient Multilingual Speech-to-Speech Translation Framework via Seamless Speech-Text Alignment and Progressive Fine-tuning
 Abstract
-Multilingual speech-to-speech translation (S2ST) aims to directly convert spoken utterances from multiple source languages into fluent and intelligible speech in a target language.
-Despite recent progress, several critical challenges persist: 1) achieving high-quality S2ST remains a significant obstacle; 2) most existing S2ST methods rely heavily on large-scale parallel speech corpora, which are difficult and resource-intensive to obtain.
-To tackle these challenges, we introduce S2ST-Omni, a novel, efficient, and scalable framework tailored for multilingual speech-to-speech translation.
-Specifically, we decompose S2ST into speech-to-text translation (S2TT) and text-to-speech synthesis (TTS).
-To enable high-quality S2TT while mitigating reliance on large-scale parallel speech corpora, we leverage powerful pretrained models: Whisper for robust audio understanding and Qwen 3.0 for advanced text comprehension.
-A lightweight speech adapter is introduced to bridge the modality gap between speech and text representations, facilitating effective utilization of pretrained multimodal knowledge.
-To ensure both translation accuracy and real-time responsiveness, we adopt a streaming speech generation model in the TTS stage, which generates the target speech in an autoregressive manner.
-Extensive experiments conducted on the CVSS benchmark demonstrate that S2ST-Omni consistently surpasses several state-of-the-art S2ST baselines in translation quality, highlighting its effectiveness and superiority.
-Index Terms—  Speech-to-speech translation, speech language model, speech-text alignment, streaming speech decoder
+Despite recent advances in multilingual speech-to-speech translation (S2ST), several critical challenges persist: 1) achieving high-quality translation remains a major hurdle, and 2) most existing methods heavily rely on large-scale parallel speech corpora, which are costly and difficult to obtain.
+To address these issues, we propose S2ST-Omni, an efficient and scalable framework for multilingual S2ST.
+Specifically, we decompose the S2ST task into speech-to-text translation (S2TT) and text-to-speech synthesis (TTS).
+For S2TT, we propose an effective speech language model that integrates the pretrained Whisper encoder for robust audio understanding and Qwen 3.0 for advanced text comprehension.
+A lightweight speech adapter is employed to bridge the modality gap between speech and text representations.
+To further facilitate the multimodal knowledge learning, a two-stage fine-tuning strategy is introduced.
+In the TTS stage, we adopt a streaming autoregressive generation approach to produce natural and fluent target speech.
+Experiments on the CVSS benchmark show that S2ST-Omni consistently outperforms existing state-of-the-art S2ST systems in translation quality, highlighting its effectiveness and superiority.
+Index Terms—  Speech-to-speech translation, speech-text alignment, speech language model, streaming speech generation
 $\dagger$ denotes the corresponding author.
 1 Introduction
-Multilingual speech-to-speech translation (S2ST) seeks to directly convert spoken input from multiple source languages into speech in a target language.
-As a critical intersection of speech technology and cross-lingual understanding, it plays a fundamental role in enabling natural and efficient multilingual spoken communication [1, 2].
-Conventional S2ST systems [3, 4, 5] predominantly adopt a statistically based cascaded architecture that decomposes the S2ST task into separate modules: automatic speech recognition (ASR) [6, 7], machine translation [8, 9], and text-to-speech synthesis (TTS) [10, 11].
-Notable examples include the Verbmobil project [5], which employs a hybrid multi-engine approach to handle multilingual business dialogues, and the corpus-based framework proposed by [4], integrating multilingual speech recognition, text translation, and speech synthesis to achieve high-quality translation among English, Japanese, and Chinese.
-Nonetheless, cascaded pipelines inherently suffer from error propagation, rigid module interfaces, and limited real-time performance, restricting their effectiveness in complex practical applications [12].
-With rapid advances in machine and deep learning, direct S2ST methods have emerged [13, 14, 12, 15], demonstrating better performance compared to traditional S2ST systems.
-TranSpeech [13] introduced a non-autoregressive (NAR) speech-to-speech translation model that addresses acoustic multimodality through a bilateral perturbation mechanism—combining style normalization and information enhancement to learn deterministic linguistic representations.
+Multilingual speech-to-speech translation (S2ST) aims to directly convert utterance from multiple source languages into a target language speech, which plays a pivotal role in cross-language communication, education, and medical care [1, 2].
+Conventional S2ST systems [3, 4, 5] normally adopt a statistically based cascaded architecture that decomposes the S2ST task into separate modules: automatic speech recognition (ASR) [6, 7], machine translation [8, 9], and text-to-speech synthesis (TTS) [10, 11].
+Notable examples include the Verbmobil project [5], which employs a hybrid multi-engine approach to handle multilingual business dialogues, and the corpus-based framework proposed by [4], integrating multilingual speech recognition, text translation, and speech synthesis to achieve translation among English, Japanese, and Chinese.
+Nonetheless, cascaded pipelines inherently suffer from error propagation, rigid module interfaces, and limited real-time performance, restricting their effectiveness in complex practical scenarios [12].
+With recent advances in deep learning, direct S2ST methods have emerged [13, 14, 12, 15], demonstrating better performance compared to previous methods.
+TranSpeech [13] introduced a non-autoregressive (NAR) speech-to-speech translation model that addresses acoustic multimodality through a bilateral perturbation mechanism that combines style normalization and information enhancement to learn deterministic linguistic representations.
 Afterwards, DASpeech [14] used a two-pass NAR architecture that first generates target text using a directed acyclic graph-based linguistic decoder and then produces target speech with a FastSpeech 2 [16] acoustic decoder, showcasing the effectiveness and potential of unified models to capture cross-lingual speech mappings.
-Despite their impressive performance, these methods still rely heavily on large-scale, strictly aligned parallel speech corpora, limiting scalability to low-resource languages.
-Moreover, there remains substantial room for improving translation accuracy.
-To overcome these problems, we propose S2ST-Omni, an end-to-end scalable speech language modeling framework for multilingual speech-to-speech translation.
-The core principle of our proposed S2ST-Omni lies in the efficient and seamless integration of large-scale multimodal pretrained models.
-This design significantly reduces reliance on parallel speech data while enabling efficient and high-quality S2ST.
-To elaborate, S2ST-Omni leverages powerful large-scale pretrained models: Whisper and for robust audio comprehension and Qwen 3.0 large language model (LLM) for deep text understanding.
-Meanwhile, a lightweight speech adapter is proposed to bridge the modality gap between speech and text, facilitating effective integration of multimodal pretrained knowledge.
+Despite impressive performance, these methods still rely heavily on large-scale aligned parallel speech corpora, limiting scalability to low-resource languages.
+Moreover, there remains substantial room to improve the translation accuracy.
+To mitigate this limitation, ComSpeech [12] introduced a compositional framework that leverages existing S2TT and TTS models, reducing dependence on triplet data.
+However, due to its reliance on a vocabulary adaptor and phoneme-level embeddings, ComSpeech generates relatively shallow phonetic representations with limited contextual integration.
+This inherent limitation reduces its capacity to model long-range dependencies, thereby constraining its effectiveness in handling semantically complex or ambiguous utterances.
 Fig. 1: Overall architecture of the proposed S2ST-Omni framework. LSA denotes the lightweight speech adapter.
-To further enhance translation quality while ensuring real-time usability, we employ a pretrained streaming speech decoder to synthesize the target utterance in a autoregressive (AR) fashion, making S2ST-Omni well-suited for practical deployment in real-world multilingual communication scenarios.
-Extensive experiments on the CVSS multilingual speech translation dataset demonstrate that S2ST-Omni achieves substantial improvements over current leading end-to-end and cascaded baselines, delivering higher translation quality alongside competitive latency and robustness, thereby illustrating its practical applicability and promising potential.
-2 Related Work
-This section briefly presents the background for the two-pass S2ST and speech language model for better legibility.
-2.1 Two-pass S2ST
-Due to the challenges of directly generating target speech [12], many research [17, 18] adopts a two-pass architecture, which decomposes the task into S2TT and TTS.
-While this paradigm benefits from intermediate textual supervision to enhance translation quality, it often relies heavily on large-scale triplet-parallel corpora (speech-text-speech), limiting its scalability in multilingual and low-resource scenarios.
-To address this limitation, ComSpeech [12] introduces a compositional framework that leverages existing S2TT and TTS models, reducing the dependence on triplet data.
-However, its reliance on a Vocabulary Adaptor, CTC decoding, and Phoneme Embedding results in a shallow representation at the phonetic level, lacking the capacity to model deep semantics or long-range contextual dependencies.
-As a consequence, it struggles to capture complex linguistic phenomena, particularly in semantically rich or ambiguous contexts.
-2.2 Speech Language Model
-Recent progressions in discrete audio representations [19, 20, 21] and self-supervised learning [22, 23] have led to powerful speech language models (SLMs), such as GSLM [24], VoxtLM [25], and VALL-E [26].
-These models excel in natural speech generation and zero-shot synthesis.
-Inspired by their impressive performance, SLMs have been applied to S2ST by decoupling translation and acoustic unit generation, improving translation quality and flexibility [27, 28].
-However, these approaches heavily depend on large-scale parallel speech corpora and struggle with effective speech-text modality fusion.
-In contrast, our proposed S2ST-Omni leverages pre-trained multimodal large-scale models with lightweight speech adapters to align multimodal speech and text representations, thereby reducing reliance on parallel triplet data while enhancing semantic understanding and enabling efficient end-to-end training.
+To solve these problems, we propose S2ST-Omni, an innovative and scalable framework for multilingual S2ST.
+Essentially, the core principle of S2ST-Omni lies in the efficient and seamless integration of large-scale multimodal pre-trained models, which can significantly reduce reliance on parallel speech data while enabling efficient and high-performance S2ST.
+To elaborate, we first present a speech language model (SLM) that leverages powerful pre-trained Whisper-large-v3 [17] encoder for robust audio comprehension and Qwen3.0 4B large language model (LLM) [18] for advanced text understanding.
+Meanwhile, a lightweight speech adapter (LSA) module is introduced to bridge the modality gap between speech and text, enabling efficient fusion and joint modeling of multimodal knowledge.
+To further enhance multimodal learning, a two-stage fine-tuning strategy is introduced.
+In the first stage, only the LSA module is updated to achieve initial convergence.
+Afterwards, both the LSA and Qwen LLM are jointly optimized during the second phase, where the LLM is fine-tuned using the parameter-efficient low-rank adaptation (LoRA) [19] method.
+Moreover, to improve translation quality and ensure low-latency speech synthesis, we adopt a pre-trained streaming speech generation approach [20] to synthesize the target utterance in an autoregressive (AR) fashion.
+Extensive experiments on the CVSS [21] corpus demonstrate that S2ST-Omni achieves substantial improvements over current state-of-the-art (SOTA) baselines, consistently delivering the best translation quality.
 2 METHODOLOGY
@@
 2.1 System Overview
-As shown in Figure 1, our proposed S2ST-Omni framework comprises four key components: a pre-trained speech encoder, a lightweight speech adapter, a pre-trained LLM, and a pre-trained streaming speech decoder.
-Specifically, to achieve high-quality S2TT, we first leverage the pretrained Whisper model in conjunction with a lightweight speech adapter to effectively extract and align speech representations.
-These features are then transformed into representations compatible with the pretrained Qwen 3.0 LLM.
+As shown in Fig.
+1, our proposed S2ST-Omni framework comprises four key components: a pre-trained speech encoder, a learnable LSA, a pre-trained LLM, and a pre-trained streaming speech generation model.
+Specifically, to achieve accurate S2TT, we first leverage the pre-trained Whisper model in conjunction with a learnable LSA module to effectively extract and align speech representations.
+These features are then transformed into representations compatible with the pre-trained Qwen3.0 LLM.
 The entire S2TT model is optimized using cross-entropy loss to ensure accurate and efficient speech understanding and recognition.
-For efficient and real-time speech synthesis, we integrate a pretrained streaming speech decoder that generates target audio in an AR manner.
-3.2 S2TT Stage: Seamless Speech-Text Alignment
-Benefiting from training on large-scale and weakly supervised multilingual corpora, the Whisper [29] architecture has demonstrated a strong generalization ability in capturing diverse acoustic and linguistic characteristics across a wide range of languages and conditions, including low-resource scenarios.
+For efficient and fast speech synthesis, we employ a pre-trained streaming speech generation model [20] that generates target audio in an AR manner.
+2.2 SLM via Seamless Speech-Text Alignment
+Benefiting from training on large-scale and weakly supervised multilingual corpora, the Whisper [17] architecture has shown a strong generalization ability in capturing diverse acoustic and linguistic characteristics across a wide range of languages and conditions, including low-resource scenarios.
 These properties make it an ideal choice for multilingual speech understanding and translation tasks.
-Consequently, we adopt Whisper-medium as the speech encoder in our S2ST-Omni workflow.
-Let $X_{s}\in\mathbb{R}^{T}$ denote the input source utterance, where $T$ is the duration of $X_{S}$, $\bm{T}\in\mathbb{R}^{L}$ represent its corresponding target transcript, where $L$ equals its text token length.
-The Whisper model first converts the raw waveform into a log-Mel spectrogram feature $F\in\mathbb{R}^{T_{1}\times D}$, where $T_{1}$ is the number of spectrogram frames and $D$ is the acoustic feature dimension, typically set to 80.
-Afterward, the log-Mel spectrogram $F$ is passed through the Whisper encoder, which consists of a stack of transformer blocks [30] comprising multi-head self-attention and feed-forward layers.
+Consequently, we adopt Whisper-large-v3 as the speech encoder in the proposed SLM.
+Let $X_{s}\in\mathbb{R}^{T}$ denotes the input source utterance, where $T$ is the duration of $X_{S}$, $\bm{T}\in\mathbb{R}^{L}$ represent its corresponding target transcript, where $L$ equals its text token length.
+The Whisper model first converts the raw waveform into a log-Mel spectrogram feature $F\in\mathbb{R}^{T_{1}\times D}$, where $T_{1}$ is the number of spectrogram frames and $D$ is the acoustic feature dimension.
+Then, the log-Mel spectrogram $F$ is passed through the Whisper encoder, which consists of a stack of transformer blocks [22] comprising multi-head self-attention and feed-forward layers.
 The encoder also incorporates positional embeddings to retain temporal information and layer normalization for stable training.
-Through this architecture, the Whisper encoder is able to extract high-level speech representations that capture both local acoustic features and long-range contextual dependencies.
-To fully leverage the powerful semantic understanding capabilities of the LLM, it is essential to bridge the representation gap between the Whisper encoder and the Qwen 3.0 LLM.
-To this end, we present a lightweight speech adapter $\mathcal{A}(\cdot)$ based on a multi-layer projection (MLP) module, which maps the encoded speech features into a compatible representation space compatible with the Qwen 3.0 LLM.
+Through this architecture, we can extract high-level speech representations that capture both local acoustic features and long-range contextual dependencies.
+Subsequently, to fully exploit the powerful semantic understanding capabilities of the LLM, it is essential to bridge the representation gap between the Whisper encoder and the Qwen 3.0 LLM.
+To this end, we present a LSA module $\mathcal{A}(\cdot)$ based on a multilayer projection (MLP) module, which maps the encoded speech features into a compatible representation space compatible with the Qwen 3.0 LLM.
 Formally, assume $\mathcal{F}_{\text{whisper}}\in\mathbb{R}^{T_{w}\times D_{w}}$ denote the encoded speech features produced by the Whisper encoder.
@@
 $\begin{split}\mathcal{F}_{\text{adapt}}&=\mathcal{H}_{4},\end{split}$ (3)
-where $\mathcal{F}_{\text{adapt}}\in\mathbb{R}^{T_{a}\times D_{a}}$, $\mathbf{W}_{i}$ and $\mathbf{b}_{i}$ are the learnable weight matrices and bias vectors of the $i$-th linear layer, respectively, and $\sigma(\cdot)$ denotes the LeakyReLU activation function with a negative slope of 0.1.
-The operator $\prod$ here represents nested functional composition.
+where $\mathcal{F}_{\text{adapt}}\in\mathbb{R}^{T_{a}\times D_{a}}$, $\mathbf{W}_{i}$ and $\mathbf{b}_{i}$ are learnable weight and bias parameters of the $i$-th linear layer, $\sigma(\cdot)$ denotes the LeakyReLU activation function with a negative slope of 0.1, and the operator $\prod$ represents a nested functional composition.
 In this way, S2ST-Omni enables efficient alignment of speech features with the language model’s input space while maintaining computational simplicity and training stability.
-Finally, the adjusted representation $\mathcal{F}_{\text{adapt}}$ is input into the Qwen 3.0 LLM for S2TT, with the guidance of the conventional cross-entropy loss:
-$\mathcal{L}_{\text{S2TT}}=-\sum_{l=1}^{L}\log P(t_{l}\mid t_{<l},\mathcal{F}_{\text{adapt}})$ (4)
-where ${t_{l}}$ denotes the $l$-th predicted target text token, and $t_{<l}$ represents the autoregressively generated text tokens preceding the $l$-th position.
-3.3 TTS Stage: Streaming Speech Decoder
-To achieve a balance between synthesis latency and translation quality, we incorporate a pretrained streaming speech decoder, CosyVoice2 [31], into the TTS component of S2ST-Omni.
+Finally, the adjusted representation $\mathcal{F}_{\text{adapt}}$ is input into the Qwen 3.0 LLM for S2TT, under the guidance of the conventional cross-entropy loss:
+$\mathcal{L}_{\text{SLM}}=-\sum_{l=1}^{L}\log P(t_{l}\mid t_{<l},\mathcal{F}_{\text{adapt}})$ (4)
+where ${t_{l}}$ is the $l$-th predicted target text token, and $t_{<l}$ is the autoregressively generated text tokens preceding the $l$-th position.
+2.3 Two-Stage Fine-tuning Strategy
+To further improve the translation performance of the proposed S2ST-Omni, we introduce an effective two-stage progressive fine-tuning strategy, which is illustrated in Fig. 1.
+In the first stage, we freeze the parameters of the pretrained Whisper encoder and the Qwen 3.0 LLM, and update merely the LSA module.
+Generally, this phase focuses on learning an effective alignment between the acoustic representations produced by Whisper and the textual embeddings of the LLM, without disrupting the pretrained knowledge encoded in the multimodal backbone components.
+Once S2ST-Omni reaches convergence in the first phase, we jointly optimize both the LSA and the Qwen LLM, where the LLM is fine-tuned using the parameter-efficient low-rank adaptation (Lora) mechanism.
+Concretely, LoRA is applied to the key projection matrices, i.e., q_proj and v_proj within the attention module, effectively reducing training overhead while maintaining performance.
+By this means, it enables the LLM to better adapt to speech-derived representations and enhances its ability to handle complex linguistic phenomena in spoken language, such as disfluencies, code-switching, and semantic ambiguities.
+2.4 Streaming Speech Generation
+To achieve a balance between synthesis latency and translation quality, we incorporate a pretrained streaming speech generation method, CosyVoice2 [20].
+To be specific, we first use a supervised speech tokenizer1 to convert the target waveform $X_{T}$ into a sequence of discrete speech tokens $Y=[y_{1},y_{2},\ldots,y_{M}]$, sampled at a rate of 25 tokens per second, where each token $y_{i}\in\{0,1,\ldots,6561\}$.
+The predicted text output from the Qwen 3.0 LLM, denoted as $\bm{\hat{T}}$, is subsequently fed into a text-to-speech language model, which autoregressively generates the predicted speech token sequence $\hat{\mathbf{Y}}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{M}]$.
+Following this, $\hat{\mathbf{Y}}$ is reconstructed into target Mel-spectrogram features $\bm{\hat{\mathbf{F}}}_{\mathrm{mel}}$ using a chunk-aware causal conditional flow matching (CFM) model [23, 24, 25], conditioned on prompt speaker embeddings.
+In particular, every $K$ generated speech tokens are grouped as a chunk, which serves as the input unit for Mel-spectrogram generation.
+Last, the target Mel–spectrograms are fed into a pretrained HifiGAN Vocoder2 to generate the converted utterance.
+In our case, we directly utilize the pretrained CFM3 and HifiGAN without any fine-tuning.
+1https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B/blob/main/speech_tokenizer_v2.onnx2https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B/blob/main/flow.pt3https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B/blob/main/hift.pt
+3 EXPERIMENTS
+In this section, we first describe the datasets used for the S2ST task.
+We then present the implementation details of the proposed S2ST-Omni model and briefly introduce the baseline systems used for comparison.
+Finally, we report and analyze the overall experimental results.
+3.1 Experimental Settings
+3.1.1 Databases
+[…]
 Table 1: ASR-BLEU score comparison among all evaluated S2ST methods.
@@
 StreamSpeech [15] 28.45 20.93 27.25 25.54
-S2ST-Omni ??? ??? ??? ???
-Specifically, we first employ a supervised speech tokenizer1 to convert the target waveform $X_{T}$ into a sequence of discrete speech tokens $Y=[y_{1},y_{2},\ldots,y_{M}]$, sampled at a rate of 25 tokens per second, where each token $y_{i}\in\{0,1,\ldots,6561\}$.
-The predicted text output from the Qwen 3.0 large language model (LLM), denoted as $\bm{\hat{T}}$, is subsequently fed into a text-to-speech language model, which autoregressively generates the predicted speech token sequence $\hat{\mathbf{Y}}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{M}]$.
-To enhance the quality of synthesized speech, the TTS model is optimized via a cross-entropy loss function defined as:
-$\mathcal{L}_{\text{TTS}}=-\sum_{i=1}^{M}\log P(y_{i}\mid\hat{y}_{<i},\bm{\hat{T}}),$ (5)
-where $P(y_{i}\mid\hat{y}_{<i},\bm{\hat{T}})$ represents the likelihood of correctly predicting the $i$-th token, conditioned on the previously generated tokens and the text embedding $\bm{\hat{T}}$ derived from the predicted text.
-Following this, the predicted speech tokens $\hat{\mathbf{Y}}$ are reconstructed into Mel-spectrogram features of the target utterance $\bm{\hat{\mathbf{F}}}_{\mathrm{mel}}$ using a chunk-aware, causal conditional flow matching (CFM) [33, 34] model.
-This model is conditioned on prompt speaker embeddings to ensure speaker-consistent speech synthesis.
-In particular, every $K$ generated speech tokens are grouped as a chunk, which serves as the input unit for Mel-spectrogram generation.
-In our case, we fine-tune the CFM model starting from the pretrained CosyVoice2-CFM checkpoint2, while retaining the original training objective to preserve model stability and synthesis fidelity.
-More implementation details and training configurations can be referred to [31].
-Ultimately, the predicted target Mel-spectrogram $\bm{\hat{\mathbf{F}}}_{\mathrm{mel}}$ is passed through a pretrained HifiGAN Vocoder3 [35], which reconstructs the final waveform to produce the final target speech.
-1https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B/blob/main/speech_tokenizer_v2.onnx2https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B/blob/main/flow.pt3https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B/blob/main/hift.pt
-4 EXPERIMENTS
-In this section, we first describe the datasets used for the S2ST task.
-We then present the implementation details of the proposed S2ST-Omni model and briefly introduce the baseline systems used for comparison.
-Finally, we report and analyze the overall experimental results.
-4.1 Experimental Settings
-4.1.1 Databases
-All experiments are conducted on the CVSS corpus [17], a public large-scale multilingual S2ST benchmark.
-The CVSS corpus consists of aligned triplets in the format of <source speech, target text, target speech>  across 21 source languages translated into English.
-[…]
+S2ST-Omni 30.20 32.09 28.02 30.10
 3.1.2 Implementation Details
-In all experiments, we employ the AdamW optimizer to train the proposed S2ST-Omni model, with an initial learning rate set to $1\times 10^{-5}$.
-The batch sizes for the speech-to-text translation (S2TT) and text-to-speech (TTS) stages are 32 and 128, respectively.
-All models are trained for 100,000 steps using the PyTorch framework on four NVIDIA A100 GPUs.
-The system prompt for the used Qwen 3.0 4B LLM is defined as: “You are an expert in speech-to-speech translation, responsible for accurately translating multilingual input speech into corresponding English text.”
-Regarding evaluation, we adopt two widely used metrics: ASR-BLEU and BLASER 2.0, to comprehensively assess the performance of all comparative methods.
-Specifically, for ASR-BLEU, the translated speech is first transcribed into text using a pretrained ASR model4.
-The resulting transcripts are then evaluated using the BLEU score [36], calculated with the SacreBLEU toolkit5, to quantify translation quality and statistical significance.
-For BLASER, we follow [15], employing the BLASER 2.0 toolkit6 to measure cross-lingual semantic similarity between source and translated utterances.
-4https://dl.fbaipublicfiles.com/fairseq/wav2vec/wav2vec_vox_960h_pl.pt5https://github.com/mjpost/sacrebleu6https://facebookresearch.github.io/stopes/docs/eval/blaser
+For all experiments, the proposed S2ST-Omni system is trained using the AdamW optimizer with an initial learning rate of $1\times 10^{-5}$ and a batch size of 32.
+Training is conducted for 100,000 steps in each of the two fine-tuning stages within the PyTorch framework, utilizing 8 NVIDIA A100 GPUs.
+The system prompt for the Qwen 3.0 4B LLM is defined as: “You are an expert in speech-to-speech translation, responsible for accurately translating multilingual input speech into corresponding English text.”
+Regarding the evaluating metric, we employ the commonly used ASR-BLEU to assess the performance of all comparative approaches.
+In practice, the translated speech is first transcribed into its corresponding transcript using a pre-trained ASR model4, after which the BLEU score [29] is computed using the SacreBLEU toolkit5 to assess translation quality and statistical significance.
+4https://dl.fbaipublicfiles.com/fairseq/wav2vec/wav2vec_vox_960h_pl.pt5https://github.com/mjpost/sacrebleu
 3.1.3 Baselines
@@
   • StreamSpeech [15]: A unified multi-task model that performs direct simultaneous S2ST by jointly learning translation and timing policies for real-time streaming input.
-Table 2: Offline S2ST performance of StreamSpeech, evaluated with BLASER 2.0.
-[…]
-BLASER 2.0 (Unsupervised)
-UnitY 0.4467 0.4473 0.5090 0.5116
-StreamSpeech 0.4486 0.4491 0.5155 0.5178
-S2ST-Omni ??? ??? ??? ???
-BLASER 2.0 (QE)
-UnitY 3.1674 3.1772 3.3020 3.3278
-StreamSpeech 3.1779 3.1872 3.3442 3.3669
-S2ST-Omni ??? ??? ??? ???
-BLASER 2.0 (Ref)
-UnitY 3.1744 3.1965 3.2213 3.2638
-StreamSpeech 3.1989 3.2200 3.3146 3.3525
-S2ST-Omni ??? ??? ??? ???
 3.2 Results and Analysis
 To validate the effectiveness of the proposed approach, we compare the translation performance of S2ST-Omni against six state-of-the-art S2ST systems.
-A comprehensive summary of the results is provided in the Table 1 and Table 2.
+A comprehensive summary of the results is provided in Table 1.
+As shown in Table 1, S2ST-Omni consistently and significantly outperforms all baselines on the CVSS benchmark across the evaluated language pairs, with over 50% relative ASR-BLEU improvements in certain cases.
+These results highlight the strong generalization capability of the proposed framework across linguistically diverse settings.
+To be concrete, in the French $\rightarrow$ English direction, our S2ST-Omni achieves an ASR-BLEU score of 30.20, surpassing the strongest baseline (StreamSpeech, 28.45).
+This substantial improvement underscores the robustness of the propose method in handling phonetic variability and prosodic features characteristic of French.
+For German $\rightarrow$ English, the S2ST-Omni attains a 28.02 ASR-BLEU score, the highest among all language pairs as well, demonstrating its effectiveness in managing syntactic complexity and compound structures inherent to German.
+In the Spanish $\rightarrow$ English setting, S2ST-Omni achieves a remarkable 32.09 ASR-BLEU score, again outperforming all baselines by a wide margin.
+This result reflects the model’s ability to capture the rich morphological patterns of Spanish and generate fluent, accurate translations.
+In addition, despite the significant differences in phonology, syntax, and morphology among French, Spanish, and German, S2ST-Omni consistently delivers strong and comparable ASR-BLEU scores, further validating its stability and robustness across diverse linguistic scenarios.
 4 CONCLUSIONS
-In this study, we introduce S2ST-Omni, a unified and efficient framework for multilingual speech-to-speech translation.
-By decomposing the S2ST task into speech-to-text translation and text-to-speech synthesis, and integrating large-scale pre-trained models such as Whisper and Qwen 3.0 with a lightweight speech adapter, S2ST-Omni achieves high-quality translation while reducing reliance on parallel speech data.
-Furthermore, the adoption of a streaming speech generation model enables low-latency speech generation suitable for real-time applications.
+In this study, we introduce an efficient and scalable framework, S2ST-Omni, for multilingual speech-to-speech translation.
+First, we propose an effective speech language model that integrates large-scale pre-trained multimodal models (Whisper and Qwen 3.0) with a lightweight speech adapter to achieve high-quality translation while reducing reliance on parallel speech data.
+Then, a two-stage progressive fine-tuning approach is presented to further improve the translation accuracy.
+Additionally, we employ a streaming speech generation model to balance the trade-off between the synthesis latency and translation quality.
 Extensive experiments on the CVSS benchmark confirm that S2ST-Omni achieves better translation quality than existing state-of-the-art systems, demonstrating its effectiveness and superiority.